# SAM2-ELNet: Label Enhancement and Automatic Annotation for Remote Sensing Segmentation

### v1

@@ IV-E2 Quantitative Analysis
To further quantify the impact of label enhancement on the model’s generalization performance, we conducted comparative experiments under identical training conditions. Specifically, we used DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as benchmark models. We initialized all models with the same pre-trained weights, applied identical hyperparameter configurations, and trained them for 20 epochs to evaluate their adaptability and generalization across different datasets.
TABLE II: Impact of Different Label Qualities on Model Generalization Performance (Cross-Dataset Transfer Experiment)
Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
MobileNet Original Label 0.8347 0.7024 0.1073
Enhanced Label 0.8941 0.7953 0.0606
ResNet50 Original Label 0.8702 0.7525 0.0976
Enhanced Label 0.9056 0.8166 0.0597
ResNet101 Original Label 0.8491 0.7192 0.0983
Enhanced Label 0.9100 0.8261 0.0554
Experimental results in Table II indicate that although overall performance declined after cross-dataset transfer, the models still exhibited strong automatic annotation capabilities, further confirming the feasibility of the proposed approach for automatic annotation in remote sensing imagery.
IV-F Ablation Study: Effect of the Edge Attention Mechanism
@@
To investigate the impact of the edge attention mechanism on remote sensing image segmentation performance, our work conducted an ablation study. By comparing model configurations with and without the EAM, our work evaluated differences in both qualitative visualization and quantitative metrics to assess its contribution to improving segmentation accuracy.
IV-F1 Qualitative Analysis
@@
Fig.8 presents the segmentation results obtained with (left) and without (right) the Edge Attention Mechanism.
Figure 8: Impact of the Edge Attention Mechanism on Segmentation Results. (b) Without Edge Attention Module, (c) With Edge Attention Mechanism; (e) Without Edge Attention Mechanism, (f) With Edge Attention Mechanism.
@@
Experimental results show that the Edge Attention Mechanism significantly improves segmentation accuracy along object boundaries. In regions with low contrast or blurred edges, this module captures finer details, generating smoother and more continuous segmentation boundaries while reducing jagged or fragmented edges. In contrast, models without the EAM perform worse in these areas, often distorting shapes and creating ambiguous boundaries, which lowers segmentation quality.
@@
Figure 9: Comparison of Different Labeling Methods. (a) Original SAR Image, (b) Decoder Output, (c) Threshold-Based Binarization, and (d) Proposed Labeling Method.
IV-F2 Quantitative Analysis
To further quantify the contribution of the EAM, our work conducted a comparative analysis under identical training conditions, evaluating model performance with and without the module on key metrics. The results are presented in Table III.
TABLE III: Impact of the Edge Attention Mechanism on Model Performance (Ablation Study)
EAM ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
92.1% 72.9% 0.701
95.2% 81.6% 0.412
As shown in Table III, integrating the Edge Attention Mechanism significantly improves both ACC and mIoU while reducing the final loss. These results demonstrate that the EAM effectively enhances local boundary precision, enabling the generation of more stable and accurate segmentation labels.
IV-G Efficiency Evaluation
To evaluate computational efficiency, we compare the training time per epoch across different deep learning segmentation models. We conduct all experiments on an NVIDIA 4090D GPU using a training dataset of 3,101 images. The results are shown in Table IV.
TABLE IV: Comparison of Single Epoch Training Time Across Different Segmentation Models (Efficiency Evaluation)
@@ IV-G Efficiency Evaluation
Experimental results indicate that our model achieves a significantly lower training time per epoch compared to other models. Specifically, it reduces training time by approximately 45.8% compared to DeepLabv3+ ResNet50 and 60% compared to DeepLabv3+ ResNet101. These findings highlight the superior training efficiency of our model, enabling faster convergence and lower computational costs.
IV-H Labeling Experiment Analysis
The model’s decoder output classifies each pixel to distinguish between the ocean and oil spills, producing a single-channel grayscale image. Each pixel value represents the probability of belonging to either the ocean or an oil spill. Unlike directly generating binary labels, this output consists of continuous grayscale values ranging from 0 to 255, providing more detailed confidence information.
During the annotation process, threshold-based segmentation inevitably leads to the loss of details, particularly affecting boundary precision and morphological accuracy. As shown in Fig.9, the proposed labeling module effectively preserves more details, especially in the edge contours of oil spill regions. Experimental results further demonstrate that directly using grayscale probability maps allows for a more refined representation of confidence levels across different regions. This approach retains richer information for subsequent analysis and decision-making, making it potentially more advantageous in scenarios requiring human-assisted judgment.
@@ V Conclusion
Our work proposes an edge-enhanced labeling network, called SAM2-ELNet, which combines a labeling module and an edge attention mechanism. This model not only effectively solves the problems of label detail loss, fragmentation, and inaccurate boundaries in manually annotated data, but also has high training efficiency and strong generalization ability. To evaluate the quality of the generated annotations, we conducted experiments on the manually annotated Deep-SAR Oil Spill (SOS) dataset and compared the performance of DeepLab series models trained with original and enhanced labels. The experimental results show that under the same experimental conditions, the model trained with enhanced labels achieves about a 5% improvement in mIoU and accuracy, while also showing faster convergence and a lower final loss value, which indicates that the improved annotations better approximate the real data distribution. For cross-dataset transfer, the model shows strong generalization ability and is able to generate high-fidelity pseudo labels for new datasets. This approach provides a feasible solution for efficient processing of large-scale remote sensing images.
Experiments on synthetic aperture radar (SAR) imagery confirm the model’s effectiveness. Future research will expand its application to multispectral and hyperspectral remote sensing data to assess its robustness across different modalities. Developing a universal segmentation model for remote sensing remains challenging. This study inspires further optimizations in adapter modules and decoder structures to enhance transfer learning, supporting more efficient and accurate image analysis in remote sensing.

### v2

@@ IV-E2 Quantitative Analysis
To fairly assess the impact of different labels on model performance, we use the same test set across all comparisons—the original test set of the Sentinel dataset. Two models are trained separately: one using the automatically generated labels and the other using the original labels. By comparing their performance on the same test set, we can effectively evaluate the quality and practicality of the automatically generated annotations. This experiment helps analyze the model’s ability to adapt to distribution shifts and confirms its potential for cross-dataset automatic annotation tasks.
To ensure consistent evaluation, we further adopt two mainstream semantic segmentation models: DeepLabv3+ (with ResNet50 as the backbone) and U-Net. All models are initialized with the same pretrained weights and trained under identical hyperparameter settings for 20 epochs. Performance on the shared test set is compared to assess model adaptability in cross-domain scenarios and validate the feasibility and effectiveness of the proposed framework for automatic annotation.
TABLE III: Performance Comparison Between Models Trained on Manual and Auto-Annotated Labels (Sentinel Test Set)
Test Set Model Train Set ACC mIoU
Original DeepLabv3+ Manual Labels 0.8577 0.7387
DeepLabv3+ Auto Labels 0.8566 0.7352
Original U-Net Manual Labels 0.8781 0.7672
U-Net Auto Labels 0.8722 0.7649
As shown in Table III and Fig. 8, the models trained with auto-annotated labels achieved performance comparable to those trained with original labels on the Sentinel dataset, further confirming the feasibility and effectiveness of the proposed framework in remote sensing applications.
IV-F Automatic Annotation on Very High-Resolution Remote Sensing Data
To further validate the applicability of the proposed automatic annotation framework in optical remote sensing scenarios, we conducted an additional set of experiments. Unlike the previous SAR-based experiments, this evaluation focuses on the CHN6-CUG Road Dataset, a high-resolution optical remote sensing dataset. This setup allows us to assess the generalization capability and practical value of our method under different sensor types and stricter annotation standards.
To simulate a realistic “low-cost few-shot annotation + large-scale automatic annotation” workflow, only the first 30% of the training data from the dataset was used to fine-tune the model. The fine-tuned model then automatically annotated the remaining training data to generate labels. We trained two independent models: one using the original labels and the other using the automatically generated labels, and evaluated their performance on a shared test set.
This experiment aims to examine whether the proposed framework can efficiently annotate large-scale data with limited high-quality supervision and provide reliable guidance for downstream model training under resource-constrained conditions.
TABLE IV: Performance Comparison Between Models Trained on Manual and Auto-Annotated Labels (CHN6-CUG Road Dataset)
Test Set Model Train Set ACC mIoU
Original DeepLabv3+ Manual Labels 0.9581 0.6865
DeepLabv3+ Auto Labels 0.9548 0.6561
Original U-Net Manual Labels 0.9556 0.6548
U-Net Auto Labels 0.9485 0.6379
As shown in Table IV and Fig. 10, models trained on automatically annotated data exhibit slightly lower performance compared to those trained on manual labels. However, the results remain practically valuable, offering a cost-effective solution for real-world applications.
IV-G Application of the Label Quality Evaluator
To evaluate the effectiveness of the proposed Label Quality Evaluator (LQE), we conducted an additional experiment on the CHN6-CUG Road dataset. Building upon earlier experiments that compared manually annotated labels and unfiltered auto-generated labels, this study introduced a third setting—auto-generated labels filtered by LQE.
As shown in Table V, models trained with unfiltered auto-generated labels exhibited a clear decline in both accuracy (ACC) and mean Intersection over Union (mIoU) compared to those trained with manual annotations. This confirms that label noise can negatively affect segmentation performance. However, LQE was applied to screen the auto-generated labels, and model performance improved significantly. In particular, mIoU showed a notable increase.
To further explore the influence of the Iterative Enhancement hyperparameter, we also tested different loop numbers in the iterative process. As shown in Table VI, we observed that the number of loops affected model performance. Considering both performance and time cost, we recommend using 2 or 3 iterations as a balanced choice for optimal results.
These results validate the utility of the Label Quality Evaluator. LQE distinguishes high-quality labels from noisy ones, allowing the model to learn from more reliable supervision. This makes LQE a valuable component in scalable annotation pipelines, especially under limited annotation resources.
TABLE V: Experimental Results Evaluating the Effectiveness of the Label Quality Evaluator (LQE)
Test Set Model Train Set ACC mIoU
Original DeepLabv3+ Manual Labels 0.9581 0.6865
DeepLabv3+ Auto Labels 0.9548 0.6561
DeepLabv3+ Auto Labels (LQE) 0.9572 0.6705
Original U-Net Manual Labels 0.9556 0.6548
U-Net Auto Labels 0.9485 0.6379
U-Net Auto Labels (LQE) 0.9532 0.6469
TABLE VI: Experimental Results of Iterative Enhancement with Different Loop Numbers
Loop Numbers ACC mIoU
0 0.9548 0.6561
1 0.9575 0.6632
2 0.9556 0.6686
3 0.9572 0.6705
4 0.9569 0.6697
IV-H Ablation Study: Effect of the EAM
@@
To investigate the impact of the EAM on remote sensing image segmentation performance, our work conducted an ablation study. By comparing model configurations with and without the EAM, our work evaluated differences in both qualitative visualization and quantitative metrics to assess its contribution to improving segmentation accuracy.
IV-H1 Qualitative Analysis
@@
Fig. 11 presents the segmentation results obtained with (left) and without (right) the EAM.
Figure 11: Impact of the EAM on Segmentation Results. (b) Without EAM, (c) With EAM; (e) Without EAM, (f) With EAM.
@@
Experimental results show that the EAM significantly improves segmentation accuracy along object boundaries. In regions with low contrast or blurred edges, this module captures finer details, generating smoother and more continuous segmentation boundaries while reducing jagged or fragmented edges. In contrast, models without the EAM perform worse in these areas, often distorting shapes and creating ambiguous boundaries, which lowers segmentation quality.
@@
IV-H2 Quantitative Analysis
To further evaluate the contribution of EAM, we constructed two different test sets by enabling and disabling EAM, and compared them with the baseline performance on TEST-HQ. The results are shown in Table VII.
TABLE VII: Impact of the EAM on Model Performance (Compared with TEST-HQ Baseline)
Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
Original Test-NoEMA 0.8126 0.0311 0.9391 0.0108
Original Test-EMA 0.7877 0.0062 0.9328 0.0045
As shown in Table VII, the test set generated with the EAM produces evaluation results that more closely align with those on the finely annotated high-quality Test-HQ dataset, compared to the version generated without EAM. This suggests that EAM contributes to the generation of higher-quality labels. By incorporating edge-aware constraints during label generation, the model better preserves fine boundary structures, leading to more precise and reliable automatic annotations.
IV-I Efficiency Evaluation
To evaluate computational efficiency, we compare the training time per epoch across different deep learning segmentation models. We conduct all experiments on an NVIDIA 4090D GPU using a training dataset of 3,101 images. The results are shown in Table VIII.
TABLE VIII: Comparison of Single Epoch Training Time Across Different Segmentation Models (Efficiency Evaluation)
@@ IV-I Efficiency Evaluation
Experimental results indicate that our model achieves significantly lower training time per epoch compared to other models. Specifically, it reduces training time by approximately 45.8% compared to DeepLabv3+ with ResNet50 and 60% compared to DeepLabv3+ with ResNet101. Moreover, the training time of our model is comparable to that of lightweight architectures such as LOGCAN++ and DeepLabv3+ with MobileNet, demonstrating that it maintains high computational efficiency while preserving segmentation accuracy.
IV-J Discussion
In practical remote sensing segmentation tasks, obtaining large-scale, high-quality pixel-level annotations is often infeasible, especially for new regions, sensors, or task types with scarce labeled data. SAM2-ELNet provides an efficient way to generate new training data from limited annotations through automatic labeling and label enhancement. It is particularly suitable for the following scenarios:
  1. 1.
    Post-processing of coarse annotations. Given existing coarse-grained labels, SAM2-ELNet can enhance boundary accuracy and structural completeness, making the labels more suitable for refined model training or evaluation.
  2. 2.
    Low-shot annotation-driven automation. When only a small amount of annotated data is available at the early stage of a project, the model can be fine-tuned with this data and then used to automatically label the entire image set, rapidly generating full-coverage labels for pretraining or iterative refinement.
  3. 3.
    Rapid cross-region deployment. When adapting a trained model to new regions or image sources, SAM2-ELNet can be quickly fine-tuned with a small number of local samples, reducing re-annotation costs and accelerating deployment.
The framework requires minimal computational resources for fine-tuning, and in some cases, its automatically generated labels can partially replace manual annotations. This makes it suitable for remote sensing projects with limited resources and high annotation costs. For targets with vague boundaries or ambiguous semantics, manual verification may still be necessary. Future enhancements to the label quality evaluator could further improve reliability.
@@ V Conclusion
This paper proposes SAM2-ELNet, a label enhancement and automatic annotation framework tailored for remote sensing under resource constraints. It freezes the SAM2 backbone and fine-tunes lightweight modules to quickly adapt to diverse segmentation tasks with limited annotations. To improve label quality, it incorporates an EAM for boundary construction and a label quality evaluator for filtering and feedback correction.
Experiments on the Deep-SAR Oil Spill and CHN6-CUG Road datasets demonstrate that SAM2-ELNet achieves both effective label enhancement and reliable automatic annotation. With only 30% of the annotated data used for fine-tuning, the model automatically labeled the entire dataset. Training on these generated labels yielded slightly lower accuracy than the model trained with the full set of original manual annotations, but still confirmed the practicality of the automatic annotation mechanism.
Future work includes extending SAM2-ELNet to multi-class and multi-modal annotation by incorporating textual semantics for vision-language fusion. The label quality evaluator may also be redesigned as a discriminator in an adversarial framework to enable self-improving learning. Research on automatic annotation is expected to provide an efficient and scalable solution for large-scale intelligent interpretation in remote sensing.